# 61A-Bot: AI homework assistance in CS1 is fast and cheap—but is it helpful?

### v1

@@ Abstract.
Chatbot interfaces for LLMs enable students to get immediate, interactive help on homework assignments, but even a thoughtfully-designed bot may not serve all pedagogical goals. In this paper, we report on the development and deployment of a GPT-4-based interactive homework assistant (“61A-Bot”) for students in a large CS1 course; over 2000 students made over 100,000 requests of our bot across two semesters. Our assistant offers one-shot, contextual feedback, through both a “Get Help” button within a popular code editor, as well as a “get feedback” feature within our command-line autograder. These triggers wrap student code in a custom prompt that supports our pedagogical goals and avoids providing solutions directly. We discuss our development process and deployment, then analyze possible impacts of our Bot on students, primarily through student feedback and how long it takes students to complete homework problems. We ask: how does access to 61A-Bot impact homework completion time and subsequent course performance? In addition to reductions in homework-related question rates in our course forum, we find substantial reductions in homework completion time. These are most pronounced for students in the percentile, with reductions of over 30 minutes, over 4 standard deviations faster than the mean in prior semesters. However, it is not clear that these effects transfer to assignment contexts where the Bot is not available: we observe speedups in some contexts, no change in others, and some assignments later in the semester even show a slowdown instead. Though we have begun to disentangle these effects, further research is needed.
@@ 1. Introduction
The recent wide availability of Large Language Models (LLMs) has given students in introductory Computer Science (CS) courses a tempting alternative to asking for help on programming assignments—and potentially waiting hours to receive it. However, while naively used LLMs do help students solve assigned problems, they typically do so by providing correct answers along with explanations, allowing students to avoid the process of developing solutions themselves and the learning associated with this process. A number of recent reports (Denny et al., 2023a, c; Finnie-Ansley et al., 2023; Hellas et al., 2023; Kazemitabaar et al., 2024; Liffiton et al., 2023; Wang et al., 2024) present more thoughtful approaches: new systems, also based on LLMs, geared towards offering guidance and assistance without providing direct solutions.
These systems succeed in providing guidance without direct solutions, and students report finding them helpful (Kazemitabaar et al., 2024; Liffiton et al., 2023)—but how are they helpful? What impact do these systems have on student performance when integrated into a traditional CS1 course? In this paper, we start to answer these questions: we first describe the assistant we developed and deployed to 2000 students in our CS1 course over the past year. Then, we report on a preliminary observational investigation of our assistant’s impact on those students, finding that our assistant does indeed help students complete homework more quickly, but does not necessarily help them learn the material better.
We begin by reporting on our experiences developing and deploying a “TA bot” homework assistant for students in the large CS1 course (CS 61A) we run at our large public university. We call our bot 61A-Bot. Students receive bot feedback in two ways: either (1) by clicking on a “Get Help” button in their code editor (see Figure 1), receiving back a pop-up notification containing advice for how to proceed on the homework assignment, given their current code; or (2) through a “get feedback” option to the existing course “autograder” command-line tool—allowing all students access to the same feedback tool, irrespective of the choice of code editor.
@@
Our bot then identifies the particular assignment the student is working on, collects their code, and wraps these in a custom prompt for GPT-4. We designed this prompt to steer GPT-4 towards feedback that mirrors how we ourselves typically approach student questions, aligned with recent work in this area (Markel and Guo, 2021): identifying whether the student understands the question, which concepts students might need reinforcement on, and whether they have a plan, and then helping students by providing conceptual, debugging, or planning support as appropriate.
We then report on:
@@
Initial findings of students’ subjective experiences based on student surveys over two semesters, and
@@
A retrospective analysis of student homework completion patterns, looking for possible impacts our bot may have had on students’ homework experiences and overall performance.
Required by our Institutional Review Board (IRB) to avoid differential access, we did not engage in a randomized control experimental design. Instead, we rely on observational data, analyzing completion-time performance for homeworks (for which students had Bot access) and similar-in-scope labs and more complex projects (for which students did not have Bot access) in our one semester of complete deployment—comparing these with historical completion-time data from previous semesters. While this approach does not necessarily allow us to draw conclusive causal inferences, the associations we find are nonetheless suggestive and call for further inquiry.
In our analysis, we observe remarkably shorter homework completion times, after the deployment of our Bot, on comparable problems from previous semesters. Suggestive of a direct causal effect is our parallel observation that completion times for comparable labs, where the Bot is never available, show no similar shortening after our deployment. Meanwhile, completion times on projects, where the Bot is also not available, are somewhat shorter after Bot deployment, but exhibit a smaller reduction than the homeworks. Finally, we also observe weak evidence of what may be over-reliance or a dependency effect: performance on bot-never-available labs shows some performance degradation, and students report that they feel as though labs take much longer than they would if the bot were available. In line with these time reductions, we also find substantial reductions in the number of online forum questions about homework problems. We report on these observations in detail in the following sections, and conclude with implications and opportunities for future inquiry.
@@ 2. Background & Related Work
Generative models such as ChatGPT,111https://chat.openai.com/ OpenAI Codex (Chen et al., 2021), DeepMind AlphaCode (Li et al., 2022), Amazon CodeWhispere222https://aws.amazon.com/codewhisperer/, and GitHub Copilot333https://copilot.github.com/ offer promising opportunities for enriching the learning experience of students. These models have already been leveraged by educators in different areas of Computing education (Finnie-Ansley et al., 2022; Denny et al., 2023c; Hellas et al., 2023; Denny et al., 2023a), where they accelerate content generation and seem to be impacting the relevant skills students gain in introductory CS courses. Researchers have studied LLMs in areas such as generating code explanations (Leinonen et al., 2023a; Becker et al., 2023; MacNeil et al., 2023b; Donlevy, 2023), providing personalized immediate feedback, enhancing programming error messages (Leinonen et al., 2023b; Wang et al., 2024), generating discussion forum responses (Mitra et al., 2024; Liu et al., 2024), and automatic creation of personalized programming exercises and tutorials (Sarsa et al., 2022; Yuan et al., 2023; Reeves et al., 2023) to enhance the comprehensiveness of course materials.
@@
However, the integration of LLMs in CS1 instruction comes with challenges. Students could become overly reliant on automation (a concern at least as old as calculators (Demana and Waits, 2000)), potentially hindering their development of critical problem-solving skills—though recent work suggests these negative effects can be avoided, at least for programming assistance (Kazemitabaar et al., 2023). Taken to an extreme, the resulting absence of human interaction could have negative effects, alongside other ethical concerns related to plagiarism and the responsible use of LLM-generated code. To maximize the benefits of LLMs while mitigating these challenges, a thoughtful and balanced approach to their incorporation into CS1 courses is essential (Denny et al., 2023b; Mirhosseini et al., 2023; MacNeil et al., 2023a).
@@
Through deployments of LLMs as intelligent tutors, students can receive immediate, personalized support and guidance, which would ideally foster a deeper understanding of coding concepts and promote self-paced learning—just as with pre-LLM Intelligent Tutoring Systems (see (Crow et al., 2018) for a review, and (Suzuki et al., 2017) for an example). The ability of LLMs to generate tailored resources, such as new, personalized tutorials and newly-generated code examples, not only expands the available learning materials but also accommodates students’ varying learning styles and preferences—though these generated materials are not always better (Pardos and Bhandari, 2023). One recent RCT of an error message enhancement system (Wang et al., 2024) begins to shed some light on how different approaches to hint generation, including GPT-4-augmented error messages, lead to different patterns of student behavior—finding that GPT-4-augmented messages can lead to persistent reductions in the number of attempts it takes students to resolve a bug in their code compared to a typical compiler or runtime error message. Educators should integrate LLMs as complementary tools, striking a balance between automation and human interaction while emphasizing the development of critical problem-solving skills and responsible coding practices, ultimately serving students better in their CS education.
@@
Researchers are also increasingly integrating LLM-based chatbots in courses (Donlevy, 2023; Wu et al., 2022) and online educational websites (Ofgang, 2023) to provide immediate personalized feedback, and in tools in supporting students’ development of programming skills (Prather et al., 2023; Finnie-Ansley et al., 2023; Cipriano and Alves, 2023). These include CodeHelp (Liffiton et al., 2023; Sheese et al., 2024) and CodeAid (Kazemitabaar et al., 2024), two systems (and deployments) that bear a number of similarities to our own—though those systems enable students to ask questions, while ours (we believe uniquely) builds on the student’s history of prior assistant hints and code changes in response. That body of work finds that students primarily seek answers to programming questions and debugging help; we focus on the latter here.
Our paper contributes to the body of educational research and pedagogical innovation, offering a preliminary evaluation of the impact of an LLM-based AI homework assistant on student performance on course assignments.
@@ 3. Design & Deployment
In designing our assistant, we chose to address one of the more challenging bottlenecks we faced in our large-scale course: the availability of tutors and other staff to help students when they get stuck on homework problems. In particular, we chose to focus primarily on the kinds of debugging assistance our staff are often asked for.
@@
3.1. Development & Prompt Challenges
Based on the literature, we expected that engineering dialogues through an LLM would be challenging in ways that would be hard to address (Zamfirescu-Pereira et al., 2023), and indeed we found this to be the case. Early on, in testing extended dialogue interactions, we found an increasing likelihood that GPT-4 would provide a direct solution as the conversation extended—validating our decision to start with a single-shot request rather than dialogue. Many initial challenges were addressable through prompt changes alone: for example, by including the instruction “Don’t assume a problem needs to use recursion unless it’s explicit” we avoided spurious suggestions that loop-based code should be rewritten using recursion.
One particularly frustrating challenge illustrates a number of the failures we observed early on: an over-eagerness to “correct” student code that was, in fact, already correct. One such homework problem asked students to fill in template code to compute , using the operators add and sub based on whether (see Fig. 1). With an early version of our prompt, GPT-4 would consistently question the correct solution, asking students to think about what should be done in the case .
Hoping to find a prompt that would prove robust to variations in course content, we initially avoided any problem-specific text beyond a statement of the problem itself. However, we ultimately added a problem-specific note for those homework problems that required extra steering—in the example above, this note contains a solution to the problem.
3.2. Deployment
@@
Students access our bot through two modalities. In one modality, students click on a “Get Help” button (see Figure 1) in the VS Code444Visual Studio Code, https://code.visualstudio.com editor toolbar to activate our 61A-Bot extension. The extension collects students’ code, makes a best guess of which homework problem the student is working on (as several problems often appear in a single file), and constructs a request. In the second modality, students run an autograder555 OK Client, https://github.com/okpy/ok-client in the command line which collects the student’s code, any errors from executing test cases, and constructs a request from these. These requests reach a server run by our instructional staff, which wraps the student data in our prompt, passes it to GPT-4, and logs the request and GPT-4’s response for further analysis.
@@
Students are informed when installing the software that, in using our assistant, all code they write will be sent to OpenAI via Microsoft Azure (Microsoft, 2024), and that they should not include any content in their code files (e.g., comments) that they do not want to share.
@@ 3.2.1. Deployment Timeline
We piloted and continued development on 61A-Bot throughout the academic year 2023-2024. In Fall 2023, we deployed an initial pilot of 61A-Bot to an experimental section of 400 students. This was followed by a full-scale deployment in week 7 for the approximately 1300 students across two sections. In conjunction with the wider deployment, we also enabled access through the autograder tool students could already run from the command line to validate their code against a set of test cases. In Spring 2024, all 900 enrolled students had access to both modalities of the bot from the start of the semester. Our qualitative analysis thus draws on both semesters, while our quantitative analysis is focused on the full deployment in Spring 2024.
4. Findings
Students’ adoption of 61A-Bot was immediate, and usage exploded once we integrated access into our autograder. In this section, we detail student assessments of their experiences, followed by an investigation of the possible impacts of 61A-Bot on homework completion times, exam score distributions, and reliance on course staff. Unless otherwise noted, statistical accounts in this section come exclusively from Spring 2024, the semester in which our fully-developed Bot was deployed for all students (see §3.2.1 for deployment details).
@@ 4.1. Usage & Reliance on Staff
Usage patterns suggest that students are finding 61A-Bot helpful, returning to it multiple times as they engage in homework: across our pilot and full deployment semesters, over 2000 students made a total of 105,689 requests of our bot. The median student in our full deployment semester made 25 requests to our bot, rising to 80 requests for the student at the 95th percentile. As expected, usage increases as the assignment deadline nears and is concentrated in the late afternoons and evenings—a pattern similar to the usage reported in (Liffiton et al., 2023; Kazemitabaar et al., 2024)—with a peak request rate of 291 requests/hour.
Over 93% of 61A-Bot queries were made via the autograder instead of VS Code, a skew likely explained by the “get feedback?” query to students on every run of the command-line autograder, which historically students run liberally. By contrast, the “Get Help” button in VS Code requires a manual installation of an extension, in addition to the use of VS Code, which our course does not require.
This engagement correlates with a reduction in help requests on our online discussion forum for students to receive asynchronous help. There is a substantial (30%!) decrease in the number of questions asked per student throughout from Spring 2023 to Spring 2024, from 1.741 to 1.185 questions per student. The impact on homework-specific questions is even larger, showing a 75% decrease from 0.344 homework questions per student in SP23 to 0.088 in SP24—see Table 1. (Note the rough consistency in question counts per student in Fall 2022 and Fall 2023.)

### v2

# 61A Bot Report: AI Assistants in CS1 Save Students Homework Time and Reduce Demands on Staff. (Now What?)

@@ Abstract.
Chatbot interfaces for LLMs enable students to get immediate, interactive help on homework assignments, but even a thoughtfully-designed bot may not serve all pedagogical goals. In this paper, we report on the development and deployment of a GPT-4-based interactive homework assistant (“61A Bot”) for students in a large CS1 course; over 2000 students made over 100,000 requests of our bot across two semesters. Our assistant offers one-shot, contextual feedback, primarily through a low-friction “get feedback” prompt within the command-line “autograder” our students already run to test their code. Our Bot wraps student code in a custom prompt that supports our pedagogical goals and avoids providing solutions directly. We discuss our deployment and then analyze the impacts of our Bot on students, primarily through student-reported feedback and tracking of student homework progress. We find reductions in homework-related question rates in our course forum, as well as substantial reductions in homework completion time when our Bot is available. For students in the percentile, these reductions typically exceed 30 minutes per assignment, over 4 standard deviations faster than the mean in prior semesters. Finally, we conclude with a discussion of these observations, the potential impacts on student learning, as well as other potential costs and benefits of AI assistance in CS1.
@@ 1. Introduction
The recent wide availability of Large Language Models (LLMs) has given students in introductory Computer Science (CS) courses a tempting alternative to asking a human TA for help on programming assignments—and potentially waiting hours to receive it. However, while naively used LLMs do help students solve assigned problems, they typically do so by providing correct answers along with explanations, allowing students to avoid the process of developing solutions themselves and the learning associated with this process. A number of recent reports (Denny et al., 2023a, b; Finnie-Ansley et al., 2023; Hellas et al., 2023; Kazemitabaar et al., 2024; Liffiton et al., 2023; Wang et al., 2024) present more thoughtful approaches: new systems, also based on LLMs, geared towards offering guidance and assistance without providing direct solutions.
Both students and instructors are reported to find these systems helpful (Kazemitabaar et al., 2024; Liffiton et al., 2023). Recent studies of deployed LLM-based assistants have included analyses of what affordances are most appreciated and what kinds of functionality students are most likely to use (Kazemitabaar et al., 2024; Denny et al., 2024); how to effectively design systems with guardrails to reduce instances of solution-sharing (Liffiton et al., 2023); and what kinds of error messages are most correlated with reductions in error rates (Wang et al., 2024). There is a growing interest in understanding the landscape of these new systems and their impacts on students learning computer science (Prather et al., 2024)—but while much is known about how these systems are designed and what seems to work well, comparatively less has been reported on how these systems impact specific courses.
@@
In this experience report, we first present our own assistant, a low-friction “61A Bot” with 24/7 availability that offers feedback on every run of an “autograder” that students use liberally to test their code-in-progress. Our Bot constructs a GPT-4 request using our custom prompt, homework question text, student code, and autograder error output (where available), returning its response to students. The prompt is itself designed to steer towards feedback that mirrors how we ourselves typically approach student questions, aligned with recent work in this area (Markel and Guo, 2021): identifying whether the student understands the question, which concepts students might need reinforcement on, and whether they have a plan, and then helping students by providing conceptual, debugging, or planning support as appropriate.
Then, we examine our Bot’s impact on our CS1 course—a large (-student) Python-based course targeted at CS majors. Ultimately, we ask: What impact does this deployment have on our course, on students and on staff? First, we find that students assess a majority of Bot-provided hints as being helpful, in line with prior work. Second, through a retrospective observational study aimed at quantifying impacts, we find that requests for homework help to our course forum drop dramatically (75%) after our Bot is deployed, and that students spend substantially less time completing their homework—25-50% less time, often more than 30 minutes faster, a 4-8 standard deviation reduction from the mean completion time for the same assignments in prior semesters.
We note that while these impacts do not necessarily imply improvements in student learning outcomes, speeding up homework completion times without increasing learning could still be a positive outcome: the time saved on completing traditional assignments could be harnessed by instructors to increase learning in other ways. In our final section, we discuss some of the costs and potential benefits of deploying LLM-based assistance in CS1, exploring possible learning impacts, reductions in the visibility of student challenges to staff, and opportunities for future research.
Thus, our main contributions are:
A rich description of our low-friction CS1 assistant and its deployment in our course;
@@
Findings of students’ subjective experiences based on student surveys over two semesters; and
@@
A retrospective, observational analysis of student homework completion patterns, identifying quantitative impacts our bot had on staff demand and students’ homework experiences.
@@ 2. Background & Related Work
Generative models such as ChatGPT,111https://chat.openai.com/ OpenAI Codex (Chen et al., 2021), Amazon CodeWhisperer222https://aws.amazon.com/codewhisperer/, and GitHub Copilot333https://copilot.github.com/ offer promising opportunities for enriching the learning experience of students. These models have already been leveraged by educators in different areas of Computing education (Finnie-Ansley et al., 2022; Denny et al., 2023b; Hellas et al., 2023; Denny et al., 2023a), where they accelerate content generation and seem to be impacting the relevant skills students gain in introductory CS courses. Researchers have studied LLMs in areas such as generating code explanations (Leinonen et al., 2023a; Becker et al., 2023; MacNeil et al., 2023b; Liu et al., 2024a), providing personalized immediate feedback (Bassner et al., 2024), enhancing programming error messages (Leinonen et al., 2023b; Wang et al., 2024), generating discussion forum responses (Mitra et al., 2024; Liu et al., 2024b), and automatic creation of personalized programming exercises and tutorials (Sarsa et al., 2022; Yuan et al., 2023; Reeves et al., 2023) to enhance the comprehensiveness of course materials.
@@
However, the integration of LLMs in CS1 instruction comes with challenges. Students could become overly reliant on automation (a concern at least as old as calculators (Demana and Waits, 2000)), potentially hindering their development of critical problem-solving skills—though recent work suggests these negative effects can be avoided, at least for programming assistance (Kazemitabaar et al., 2023). Taken to an extreme, the resulting absence of human interaction could have negative effects, alongside other ethical concerns related to plagiarism and the responsible use of LLM-generated code. To maximize the benefits of LLMs while mitigating these challenges, a thoughtful and balanced approach to their incorporation into CS1 courses is essential (Mirhosseini et al., 2023; MacNeil et al., 2023a; Denny et al., 2024).
@@
Through deployments of LLMs as intelligent tutors, students can receive immediate, personalized support and guidance, which would ideally foster a deeper understanding of coding concepts and promote self-paced learning—just as with pre-LLM Intelligent Tutoring Systems (see (Crow et al., 2018) for a review, and (Suzuki et al., 2017) for an example). The ability of LLMs to generate tailored resources, such as new, personalized tutorials and newly-generated code examples, not only expands the available learning materials but also accommodates students’ varying learning preferences—though these generated materials are not always better (Pardos and Bhandari, 2023). Educators should integrate LLMs as complementary tools, striking a balance between automation and human interaction while emphasizing the development of critical problem-solving skills and responsible coding practices, ultimately serving students better in their CS education.
@@
Researchers are also increasingly integrating LLM-based chatbots in courses (Liu et al., 2024a; Wu et al., 2022) and online educational websites (Ofgang, 2023) to provide immediate personalized feedback, and in tools in supporting students’ development of programming skills (Prather et al., 2023; Finnie-Ansley et al., 2023; Cipriano and Alves, 2023). These include CodeHelp (Liffiton et al., 2023; Sheese et al., 2024) and CodeAid (Kazemitabaar et al., 2024), two systems (and deployments) that bear a number of similarities to our own—though those systems enable students to ask questions, while ours (we believe uniquely) integrates feedback directly into the tool students already use to execute that code, and then builds on the student’s history of prior assistant hints and code changes in response.
@@ 3. Design & Deployment
Our deployment focused on providing students help with homework problems in part to address frequent student feedback from prior terms about long wait times for TA support for homework problems. In particular, we chose to focus primarily on the kinds of debugging assistance our staff are often asked for.
@@
3.1. Course Details & Deployment
Our course covers most of the typical CS1 content, plus a few additional topics, and we report here on the CS1 portion of our course. These modules are taught using Python, to a student population made up predominantly of CS majors and intended majors ( majors/intended), with substantial prior preparation ( any prior CS course, including high school). These demographics are broadly consistent across the semesters we examine historically in this report, and variations do not correlate with the results described here. (Fall and Spring semester students do differ demographically due to how course requirements are structured at our institution, and so we treat those as separate populations for this work.) In our course, content is typically introduced in lecture, and then reinforced in three assignment types: first, “labs,” which include a mixture of notional machine reinforcement (e.g., tracing, “What would Python do?”) and traditional homework problems, completed in small groups without access to 61A Bot. These are followed by “homework,” which students complete individually with access to 61A Bot. Finally, students complete “projects” that bring together multiple concepts towards a single goal, over a few weeks, again without access to 61A Bot. In addition, students were prohibited by course policy from using ChatGPT or other similar AI-based systems for help across all assignment types—but we do not have the means to enforce this.
3.1.1. Using 61A Bot
Students primarily (93%) access our Bot through an autograder444 OK Client, https://github.com/okpy/ok-client in the command line which provides the result of running the student’s code on a set of test cases; if any fail, the student is then asked whether they would like to receive feedback on their code from 61A Bot (see Figure 1). If they do, the autograder collects the student’s code, any errors from executing test cases, and constructs a request from these.
@@ 3.1.1. Using 61A Bot
Students using VS Code555Visual Studio Code, https://code.visualstudio.com (our course does not prescribe a specific editor) can install an extension to enable a “Get Help” button in the toolbar. When activated, the extension collects students’ code, makes a best guess of which homework problem the student is working on (as several problems often appear in a single file), and constructs a request.
In both cases, the collected code includes the student’s full assignment code file for that assignment, which includes code for multiple questions. Requests thus include: our prompt, along with any problem-specific notes; the text of the specific homework problem the student is requesting help for (typically one of 4-6 such problems within the assignment); up to 3 pairs of prior-code/Bot-response text (if the student has made prior requests); the student’s current code (for all problems within the assignment); and, finally, any error text from failed test cases. These requests are sent to a server run by our instructional staff, which passes it to GPT-4 and logs the request and GPT-4’s response for further analysis.
@@
Students are informed when installing the software that, in using our assistant, all code they write will be sent to OpenAI via Microsoft Azure666https://ai.azure.com/, and that they should not include any content in their code files (e.g., comments) that they do not want to share.
@@ 3.1.2. Deployment Timeline
We piloted and continued development on 61A Bot throughout the academic year 2023-2024. In Fall 2023, we deployed an initial pilot of 61A Bot to a section of 400 students. This was followed by a full-scale deployment in week 10 (after the CS1 portion of our course) for the approximately 1400 students across both sections of the course. In conjunction with the wider deployment, we also enabled access through the autograder tool students could already run from the command line to validate their code against a set of test cases. In Spring 2024, all 900 enrolled students had access to both modalities of the bot from the start of the semester. Our qualitative analysis thus draws on both semesters, while our quantitative analysis is focused on the full deployment in Spring 2024 where we can more confidently compare like-for-like with a full-term deployment and historically comparable student populations.
4. Outcomes
Students’ adoption of 61A Bot was immediate, and usage exploded once we integrated access into our autograder. In this section, we detail usage patterns and changes in reliance on course staff, report on student assessments of their experiences, and finally offer an investigation of the possible impacts of 61A Bot on homework completion times. Unless otherwise noted, statistical accounts in this section come exclusively from Spring 2024, the semester in which our fully-developed Bot was deployed for all students (see §3.1.2 for deployment details).
@@ 4.1. Usage & Reliance on Staff
Usage patterns show that students are returning to 61A Bot multiple times as they engage in homework: across our pilot and full deployment semesters, over 2000 students made a total of 105,689 requests of our bot. The median student in our full deployment semester made 25 requests to our bot, rising to 80 requests for the student at the 95th percentile. As expected, usage increases as the assignment deadline nears and is concentrated in the late afternoons and evenings—a pattern similar to the usage reported in (Liffiton et al., 2023; Kazemitabaar et al., 2024)—with a peak request rate of 291 requests/hour.
This engagement correlates with a reduction in help requests on our online discussion forum for students to receive asynchronous help. There is a substantial (30%) decrease in the number of questions asked (scaled to total enrollment) throughout the semester from Spring 2023 to Spring 2024, from 1741 to 1185 questions per thousand students. The impact on homework-specific questions is even larger, showing a 75% decrease from 344 homework questions per thousand students in Spring ’23 to 88 in ’24—see Table 1. (We include scaled question counts for Fall 2022 and Fall 2023 as data points illustrating the level of consistency across semesters.)